Jamwilli3:

A baby gains 11 pounds in its first year of life. The baby gained mc013-1.jpg pounds during the first four months and mc013-2.jpgpounds in its second four months. How much did the baby gain in the last four months?

Gdeinward:

Can you please post the question a little better? It is a little confusing

lowkey:

please upload the corresponding image to this question.

Mercury:

full question: A baby gains 11 pounds in its first year of life. The baby gained 4 1/4 pounds during the first four months and 3 1/2 pounds in its second four months. How much did the baby gain in the last four months? since (first four months) + (second four months) + (last four months) = 1 year, simply subtract 1 year - (first four month growth) - (second four month growth) to get last four months or in other words, 11 - (4 1/4) - (3 1/2) to do this by hand, you could convert the fractions to decimals (ex: 4 1/4 = 4.25) or improper fractions (ex: 4 1/4 = 17/4)